7. How many signals can be made with 5 different flags by raising them any number at a time?
Show full solution, thank you!

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


The flags could either be raised or lowered. We have two options per flag, and 5 flags total.

This gives 2^5 = 32 different configurations.

You can think of it like a binary number of 0s and 1s. Zero could mean "lowered" and 1 could mean "raised".
Or you can think of it like a set of light switches you can flick on or off.

Answer:32

Answer by ikleyn(52776)

.
How many signals can be made with 5 different flags by raising them any number at a time?
Show full solution, thank you!
~~~~~~~~~~~~~~~~


            As worded, this problem admits different readings, interpretations, solutions and answers.

            See my interpretation, different from that of the other tutor.

            I will assume that  5  flags are  NOT  ONLY  DIFFERENT  (they always are different),  but distinguished.
            More concretely and for clarity,  I will assume that the flags have different colors.


By raising one flag at a time, I can make 5 different signals.


By raising two flags at a time, I can make 5*4 = 20 different signals 
    (if to read them from top to bottom, or from left to right).


By raising three flags at a time, I can make 5*4*3 = 60 different signals 
    (if to read them from top to bottom, or from left to right).


By raising four flags at a time, I can make 5*4*3*2 = 120 different signals 
    (if to read them from top to bottom, or from left to right).


By raising five flags at a time, I can make 5*4*3*2*1 = 120 different signals 
    (if to read them from top to bottom, or from left to right).


In all, I can make  5 + 20 + 60 + 120 + 120 = 325 different signals,
    by raising different number of distinguishable flags at a time.

Solved  (differently).